NOAA SATOPS Morning Report: Wednesday, February 06, 2013


SATOPS Morning Report: Wednesday, February 06, 2013

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 037

GOES
All GOES operations were nominal over the past 48 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 48 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 48 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
Note: Due to failed supports at the Usingen and FCDA ground stations Jason-2 data latency requirements were exceeded. At 17:15z the Jason-2 command and control system was failed over to the backup string due to network connectivity problems with Usingen. A failed Jason-2 h/w component at FCDA was replaced at approximately 03:00z.

All other JASON-2 operations were nominal over the past 48 hours.
No anomalies affecting product delivery or s/c health & safety.

SNPP
All SNPP operations were nominal over the past 48 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
ESPC did not receive the following GINI/AWIPS images: EAST - 0915Z , 0930Zc 1145Z, 1215Z, 1230Z and 1245Z; on the WEST 0900Z, 0930Z, 1200Z, 1230Z AND 1245Z.

OMI data will not be available from J036/1600Z to J038/1200Z due to a scheduled power outage on the NRT system at NASA.

ESPC has not received HRPT data from Monterey since February 4, 2013 2327z. Problem under investigation.

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.